A gas is enclosed in a cylinder fitted with a light frictionless piston and

maintained at atmospheric pressure. When 1250 kcal of heat is added to the gas, the

volume is observed to increase slowly from 12.0 m 3 to 18.2 m 3 . Calculate:

a) the work done by the gas, and [Clue = what type of work is this?]

b) the change in internal energy of the gas. [Clue = as with most questions in this

tutorial, use up the 1 st law of ThermoD, ∆U = Q – W]

Note. 1 kcal = 4.186 kJ


1
Expert's answer
2022-01-17T10:00:04-0500

a)

"W=p\Delta V=p(V_2-V_1)=628~kJ,"

b)

"\Delta U=Q-W=4.6~MJ."